OpenStudy (anonymous):

show that two straight lines through the origin which make angles of 45 degree with straight line lx+my+n=0 are given by (l^2-m^2)(x^2-y^2)+4lmxy=0

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

Notice that the lines thru origin must intersect at 90 degrees. say the lines are : \(\large y = kx\) \(\large y = \frac{-1}{k}x\)

ganeshie8 (ganeshie8):

combining both gives the pair of lines : \(\large (y-kx)(y + \frac{1}{k}x) = 0\)

ganeshie8 (ganeshie8):

you can find the value of \(k\) by using the fact that these lines make 45 degrees angle wid \(lx + my + n = 0\) : \(\large 1 = \frac{k - \frac{-l}{m}}{1 + k(\frac{-l}{m})}\) \(\large k = \frac{m-l}{m+l}\)

ganeshie8 (ganeshie8):

plug this value in the pair of lines equation and simplify

OpenStudy (anonymous):

thanks.....